OpenStudy (anonymous):

The perimeter of a basketball court is 102 meters and the length is 6 meters longer than twice the width. What are the length and width?

OpenStudy (anonymous):

let width be X then lenth is 2x+6 perimeter=2(length+width) 102=2(2x+6+x) 51=3x+6 3x=51-6 3x=45 x=15m width=x=15m length=2x+6=2(15)+6=36m

OpenStudy (anonymous):

i verified Mehreen is right

OpenStudy (anonymous):

Thank You both very much!!!!! I will be using this site a lot more often!!!!

OpenStudy (anonymous):

welcum........